Right Triangle
Given a number N, write a program to print a triangular pattern of N lines with numbers as shown below.
Input

The input will be a single line containing a positive integer (N).
Output

The output should be N rows with numbers.
Note: There is no space between the numbers.
Explanation

For example, if the given number of rows is 4,
your code should print the following pattern
1
121
12321
1234321

Sample Input 1
4
Sample Output 1
1
121
12321
1234321

Sample Input 2
9
Sample Output 2
1
121
12321
1234321
123454321
12345654321
1234567654321
123456787654321
12345678987654321
1
Expert's answer
2021-03-05T05:14:40-0500
def triang(n):
    '''
    >>> triang(4)
    1
    121
    12321
    1234321
    >>> triang(9)
    1
    121
    12321
    1234321
    123454321
    12345654321
    1234567654321
    123456787654321
    12345678987654321
    '''
    pfx = ''
    sfx = ''
    for i in range(1,n+1):
        s = str(i)
        print(pfx + s + sfx)
        sfx = s + sfx
        pfx = pfx + s

# Program
if __name__ == '__main__':
    n = int(input('N? '))
    triang(n)